We put the preamp output directly into a multimeter and observed the same fluctuating behavior as the DAC channel was changed.
We're bypassing the relay to see if that makes any difference. The old relay wiring (to be bypassed) is shown in the attached diagram. That didn't do anything.
We're looking at filtering the DC output by 5kHz to see if there are any resonances at higher frequencies that might go away. Changing SR560 output for AC path to DC and setting gain to 1 on that unit. Also changing gain in FM31 filter bank from 1E-3 to 1. The results are shown in the attached time series. The channels FM30 and FM31 see the same thing. The only difference is that FM31 goes through an SR560 with a 0.03Hz pole (6dB).
Success by bypassing the DAC bias voltage. We switched to a 300mV bias voltage from a function generator. Doing that removed the causal PD voltage drift induced by changing the laser diode current set voltage (see the last time series). So the issue is some weird coupling into the DAC bias voltage.
[Aidan, Radhika, Nina]
We noticed that the DC channel readout (FM30) of the JPL A1 photodiode is drifting around. What we observe with no light on the photodiode, is the DC output drifiting around. It gets particularly bad when we apply voltage to other DAC channels.
For example, the attached plot shows the DC voltage from the photodiode as I change the set voltage to the laser diode driver. To be absolutely clear, the laser driver itself was completely powered off. I'm just varying the voltage going into the set point BNC connector on the back of it.
For reference, the set up is:
DAC (300mV bias) > relay > PD > relay > FEMTO preamp (1000x gain) > ADC channel FM30
QIL Cryo vacuum chamber cooldown was not as successful under the new configuration (radiation shielded by cylindrical outer + inner shields, cold finger thermally strapped to baseplate).
--> Karthik's Si cantilever workpiece was stable at 240 K.
--> Cold Finger was stable at 200 K - there is significant thermal loss between the cold finger and the workpiece.
--> Inner shield was stable at 250 K - seems to be somewhat decoupled from the baseplate; not very satisfied with the current state of the shielding.
Will need to re-examine some of the connections, which were not optimal (especially the improvised dog clamped strap-baseplate interface). Fabricating an adapter piece for the thermal strap which will be bolted 4x on a 2" x 2" grid. Might also look into a new thermal strap which could interface with baseplate directly.
Also will need to consider options to decouple outer shield from inner, and double check that shield orientation has no other solution (hoping there's an answer to the question, why would outer shield be coupled to baseplate?)
Data - cooldown 20210408 (CSV = raw, XLSX = Stephen's plots) in Box Folder [Voyager\MarinerBox\CryoEngineering\CSVlogs]
Description - 6 day cooldown. Layout described in QIL/2552. The radiation shields were installed and thermal strap was connected to baseplate. The cryocooler was turned on/off at the start/end of the data collection, and the in-vac heater was not powered on at all.
I posted a video tutorial of the diode replacement.
Aidan and I removed the old PD from the cryo chamber in order to start testing C3 (plan for tomorrow, 04/02).
- Brought chamber up to room pressure, disconnected readout wires and vacuum pump.
- Picked up chamber and placed it upside down on makeshift support stand (see pics).
- Unscrewed mounting plate and 2 inner insulation plates to reveal mounted PD.
- Had trouble unscrewing PD mount, since the screws were very close to the PD and we had to be careful not to slip and cause damage. Started with 2 side screws, then bottom (hardest), then top.
- Successfully removed PD and put away. Placed chamber components back in place without bolting in.
- Plan is to mount PD C3 in chamber tomorrow and begin testing.
To aid in taking photos of these diodes, I put a USB microscope on Anchal's desk - you can grab it from there. I use it with mac Photo Booth, but it should be easy to use with any camera application.
Also, I recommend buying a macro lens(es) for cell phones from Amazon or B&H. Label them with the QIL lab sticker so they don't disappear.
Karthik had completed in-chamber alignment efforts during a prior visit. In air alignment also completed following viewport move.
0) Removed lid for access to chamber.
--> posted demo video to ligo.wbridge QIL Cryostat HowTo Playlist.
1) Mounted RTDs to final positions - locations are Heater (cryo varnish+cigarette paper, pictured in IMG_8558 curing under weight of upsidedown bolt), Inner Shield (cryo varnish+cigarette paper, pictured in IMG_8559), Cold Finger (spring clamp), and Workpiece (spring clamp).
--> Final chamber layout may be viewed in IMG_8562
--> Note that Karthik's Si cantilever, mounted vertically in the right of the image, is NOT bolted down to the baseplate (just located on baseplate by dog clamps, held down via gravity). This will need to be investigated to enable workpiece cooling.
2) Installed radiation shield lids - no bolts to expedite the process and to see if there is any bulk motion during pumpdown and thermal cycling.
--> note that the lid for the outer radiation shield seems to interface with the current shield orientation perfectly; if there was a mismatch, it would point toward the inverted orientation being intended, but this seemed pretty definitive.
3) Installed the cryostat lid - final positioning and alignment made easier by teflon rails!
--> posted demo video to ligo.wbridge QIL Cryostat HowTo Playlist.
4) Pumped down - single button press to turn on pumping station.
--> note that it took about 1 hour for both gauges to reach a few mTorr.
5) Confirmed function of heater - set PID setpoint to 350 K and enabled outputs, observed temperature rise in heater RTD.
--> note that PID autotuning should be done at steady state with workpiece RTD, before enabling outputs again!
6) Turned on cryocooler - flip power lever and turn on green system switch.
--> start time was 10 am.
7) Started temperature datalogging to USB - press dull red indicator dot on upper right corner of CTC-100 once, and note that indicator is now bright red.
8) Remaining photos posted to the ligo.wbridge QIL Cryostat Photo Album
I'm attaching my rough first draft of the QIL photodiode testing schematic. Please provide comments for fixes/improvement!
1. Radiation Shields located (in TCS lab), unwrapped, fitted up.
Location in TCS Lab - IMG_8351
Removed lid and placed adjacent to chamber (cleared a little space, used 3 plastic flange covers to make nonmarring surface safe for lid and o-ring - IMG_8353
Fit up as installed - IMG_8354
Comments on the fit up - I looked at all of the apparent sources for insights into Rahul's original design intent - QIL elog 2276, DCC T1800308-v1, wiki for Cryo Vacuum Chamber. It appears that Rahul never decoupled the upper outer radiation shield from the cold plate, which seems like a strange omission. Chris and Raymond also appear to have been wrapping their heads around the intended layout, they came up with the fit up in QIL elog 2429 and sketch from QIL elog 2430. I will revisit their sketch at a future opportunity, but I went with something closer to 2429 as I was concerned about the height misalignments they described. Note that the height misalignment appears in Rahul's T1800308 CAD (see T1800308-v1 screenshot) so who knows what's "correct". I'll work on finalizing D2100310 CAD with radiation shield to capture the true current dimensions and fit up, to hopefully avoid such issues in the future.
2. Radiation Shields installed in Cryostat. Sequence was important here, as were a couple of improvised solutions to shortcomings of the existing parts.
Dog Clamps placed on bottom plate (to stand off bottom radiation shield bottom lid; not pictured, I think I placed some alumina washers on the dog clamps as well, not sure though anymore!). Also pictured are the usual PEEK legs for cold plate - IMG_8355
Bottom radiation shield bottom lid placed on dog clamps spacer, and bottom radiation shield cylinder placed on bottom lid - IMG_8356. Seems likely that the bottom radiation shield would be better configured upside-down.
Bolted cold plate down onto legs, with cold plate decoupled from bottom radiation shield - IMG_8357
Outer radiation shield installed and inner radiation shield installed (both needed to be tipped into place gingerly, but both cleared the cold finger cylinder with the flange removed. The heater also passed through the apertures successfully - IMG_8358
2x Alumina washers placed under outer radiation shield, inner radiation shield on cold plate - IMG_8374
Cold Flange reinstalled, though one of the brass SHCS was sheared - this was due to over torque, with 20 in*lb applied by mistake. Correct torque is 10 in*lb. The remaining 3 bolts were tightened to 10 in*lb. - IMG_8360
Top view of radiation shield apertures and cold plate grid - IMG_8375
Thermal strap interface to cold plate - dog clamps required due to strange spacing of clearance holes. - IMG_8376
- Unscrewed outer and inner insulation plates.
1. Viewport swap to nozzle that is not occluded by cryo shield = complete. All bolts on both Active Ion Gauge and Viewport have been torqued gradually (about a half turn at a time, around the clock dial) until the conflat seal was metal-to-metal. Periscope on damped optical rod was rotated to make room for replacement.
Before viewport swap - IMG_8487
After viewport swap - IMG_8502
2. Cryo RTD repair = complete. Two RTDs had been damaged during prior mounting efforts by me. I was able to repair the clamped RTD at the single damaged solder joint. I was able to repair the former Al-Block RTD by replacing the RTD element, and making a new direct attachment (not preloaded, not varnished to the aluminum block anymore)
Materials and set-up for solder repair - IMG_8491
Repaired Clamp-2 RTD - IMG_8494
Damaged Al-Block RTD - IMG_8492 (note short length between kapton strain relief and aluminum block was not ideal, one lead had already fractured and the second soon followed at the slightest touch)
Repaired, remounted Al-Block RTD - IMG_8495 (heater sandwiched underneath threaded adapter, clamp threaded into adapter, sandwiching RTD at top plane)
Remounted Clamp-2 RTD - IMG_8496 (RTD clamped at cold flange, strap is mounted)
Remaining Clamp-1 and Varnish RTDs are free - IMG_8497
Current readouts of CTC-100 controller, with repaired RTDs now behaving (note need to rename the Al-Block RTD) - IMG_8501
3. Next steps:
- Karthik to install clamps, align in-air relay, and confirm positition of radiation shield aperture.
- Remaining free RTDs to be mounted; current RTDs are mounted at Heater and Cold Flange, would be good to mount RTD at Work Piece/Clamp and at Outer Radiation Shield.
- Radiation shield lids to be installed (might be easiest to install Outer Radiation Shield RTD after installing lid)
- Mount lid, install bolts, pump down, turn on cryo cooler, the usual!
[Aidan, Jon, Chris W, Ian]
Summary: We rebuilt the Cymacs C4TST today to get FM31_OUT into frames
1982 sudo /sbin/rmmod c4tst c4iop
1983 cd /opt/rtcds/caltech/c4/target/c4iop/scripts/
1985 cd ../../c4tst/scripts/
1987 systemctl start email@example.com
1988 cd .././../
1990 cd gds
1992 cd awgtpman_startup/
1996 systemctl restart firstname.lastname@example.org
1998 systemctl status email@example.com
1999 systemctl stop firstname.lastname@example.org
2000 sudo systemctl restart email@example.com
Added Simulink > Model-Wide Utilities > Model Info block to c4tst.mdl. Text inside that block is:
Now following https://nodus.ligo.caltech.edu:8081/QIL/2336
And it failed. See attached screenshot. Then I copied c4tst.mdl to the simLink directory. Compile still failed.
Noticed that the DAC channels were not producing a corresponding output in the real world (I changed the Laser Current FM12 value and got not corresponding change on the laser diode driver display).
Sent the following to Chris: "Can you log into the QIL FB4 workstation to see if there is an issue with the DAC? I restarted the C4TST model last week and I don’t seem to have working DAC outputs anymore. The ADC channels still work and the model appears to be running. It just seems that I can’t output any voltages."
After observing that the "DK" (DACKILL) bit in the state word on the IOP status screen was red, the resolution to this was to restart the IOP and TST models.
Adding fb4:/usr/share/advligorts to QIL-WS2 to /etc/fstab file
Should help access to CDS_PARTS model file in Simulink on QIL-WS2
Except access is denied by FB4
MATLAB license had expired on QIL-WS2 so I had to activate it again.
entered just before (Wed Mar 17 16:06:37 2021) to borrow a mini-circuits filter (SLP-100)
I added a 7 minute video to the DCC that shows how to operate the HiCube 80 Eco pumping station.
Also added a "Tutorial video" category to the elog.
I pumped the chamber down and added LN2 today. The pressure was slowly rising - it was about 20m Torr in the chamber when I added the LN2. Per Raymond's instructions, I added about a third of a container of LN2. This got the temperature down to about 89K (when I had 20W running in the heater). It stayed there for about 25-30 minutes.
I turned off the heater and left the LN2 to boil off. You could see the cloud coming out of the top (the plume height would increase proportionally to the heat in the heater).
Eventually the LN2 evaporated and the shield temperature started to increase back to room temperature. As of this post it is 282K (which took about 5 hours).
The PD thermistor is not currently registering. However, the temperature of the PD can be inferred from the shield temperature (see aLOG 2517).
The rate of increase in temperature was much faster than the previous test - see second time series. I wonder if the thermal mass of the shields in the Feb 2020 test was cooled down a lot more due to 5 hours at 80K in that test - thus reducing the overall ambient load on the inner shield ...
I pumped the small vacuum volume down but the pressure started rising as soon as I turned off the vacuum pump. Closing the main valve to the pump and the valve to the chamber did little to change the leak rate. So the main leak seems to be from the volume around the pressure gauge - best guess, the section and O-ring that I connected to the chamber yesterday.
Vacuum pressure was recorded from vacuum gauge to text file in Python (using pyserial). Haven't got this into EPICS just yet.
Link to ligo.wbridge QIL Cryostat HowTo Playlist Cryostat on youtube - not super user-friendly as of yet, but populated with a couple of videos so far.
Link to ligo.wbridge QIL Cryostat Photo Album on google photos - not well curated, currently just a dump.
I recorded a 15 minute overview that describes the JPL PD set up and how to operate it. I'm in the process of embellishing the operation procedure (previous version can be found here: eLOG 2476).
I measured the power incident on the cryo chamber viewport and the reference PD reading to calibrate the incident power. Data is attached. Power meter head = S148C.
I ran the bright PD test on the photodiode currenlty in the vacuum chamber. The test was run at air and room temperature. I aligned the 2um laser onto the PD using the piezo mirror and the readout from the preamp. I then switched to the Keithley and ran the bright scan with the "runsweep.py" script. I actually ran the scan at multiple laser diode current settings by varying the control voltage into the diode driver. The change in response wrt control voltage looks linear but I need to run an analysis on it.
The data is stored in /home/controls/JPL_PD/data/20210303_bright_scans
is the x-axis in units of seconds? I think if we are clever, we should be able to look at a couple of the thermal time constants and figure out where the heat leaks are.
[updated with reference to data set, cleaner plot, images of chamber configuration]
Data - cooldown 20210205 (CSV = raw, XLSX = Stephen's plots) in Box Folder [Voyager\MarinerBox\CryoEngineering\CSVlogs]
Description - 5.5 hour cooldown with data, which was then allowed to continue for a total of 96 hours (but data collection failed for the long stretch, except for a snapshot of the final state). The cold flange was below 100 K after 6 hours, and leveled off at about 80 K. The vacuum pressure was steady at 3 microTorr throughout, via a roughing line connecting the external and internal volumes (mitigated losses in external connection volume, with no areas dramatically cold to touch). The cold flange was radiating to room temperature surroundings, as the radiation shields were not installed. The cryocooler was turned on/off at the start/end of the data collection, and the in-vac heater was not powered on at all.
StephenA, RaymondR remotely assisting (off payroll haha)
It seems that we won't likely receive the intended hand-off resources (especially of note is that Raymond can't seem to find the videos he made, wherein he guides through operations of the vacuum system and cryocooler). Raymond has been kind enough to support via Zoom as needed so that things can progress with some sort of guidance.
I'll stay on top of the lessons learned and dump these, along with photos and other resources in the log. I'll also make weekly visits with the intent of making continued progress.
1) What is the current state of the QIL Chamber?
- Raymond left the vacuum line shorting the "external volume" with the cross, cryocooler, etc. directly to the "main volume" because the losses between the "external volume" and the feedthrough entering the chamber were too large. The system was down in the e-5 torr range. Ref IMG_8019, with flex hose connecting bottom of 4-way cross to side of chamber.
- To implement this vacuum arrangement, there was a key component put aside on the table - IMG_8024 is a T which connects the roughing line to both the external volume and the main volume, using the flex line from IMG_8019 and the components pictured in IMG_8020.
- The copper feedthrough has all clamps attached, such that temperature measurements are being made on the adapter copper rod, which has a bolt pattern for thermal straps. Sensor names reflect current locations.
- We inspected everything (cryocooler connections, vacuum gauges, temperature logging), and pronounced it "ready to go" at the end of my work day.
2) What did I learn today?
- Cryocooler has only one setting, and temperature control must be engineered at the output using thermal contact, emissivity, etc.
- Formatting of USB is the main error that can befall the CTC100 datalogger. If the red dot in the upper right corner of the screen does not light up bright when it is tapped (this starts datalogging), then there is something wrong. Easy enough to test by removing the USB when the red dot is dim (datalogging paused) and checking whether there are log contents.
- Raymond's focus with the QIL chamber had been on answering the question, "can we cool down the cryocooler's connection (copper linkage which passes into the chamber) adequately?" He had never successfully obtained a cooldown that was below 150 K, and the primary limitation appeared to be related to high pressures in the "external volume".
3) What are we up to next?
- The next time I come in, I will be turning on the cryocooler and datalogging first thing, and I will hopefully have cooldown trends to share in the log.
- If those trends are > 150 K, I was advised that the next thing to do would be to bring the "external volume" out of the equation, and directly attach the cryocooler to the copper feedthrough linkage. This would be one way to demonstrate the least-lossy, best case scenario.
- If < 150 K, I am told that Karthik may be ready to move in for some measurements. If not, I would be interested in dropping in the suspended, shielded Silicon dummy (currently standing by) and seeing if we can measure a successful (< 150 K) cooldown on the Si mass.
4) Can we increase the height of the chamber?
I've shared lots of images related to the question of extending the height of the chamber. Here are my thoughts:
- Raw measurements - Ceiling = 0", Crane ~ -12", Crane hook ~ -16", Chamber Lid ~ -26", Chamber Base ~ -40", Table ~ -43".
--> Not much space above the surface of lid, currently about 10" of range for a possible extension.
--> Actual useable range is less, due to real world limitations such as the height of lifting straps, interference with the angled crane arm, etc.
- It would require a clever solution to increase the crane height (spacer at base? extended height model?) or lower the lid height wrt the crane (position on lower table? lower the table on its leveling feet?) to buy a few more inches.
- Current allowed object height ~8" (could be extended to about 10" with modified PEEK spacers at base); would we benefit greatly from having a ~16" allowed object height? Or do we need to get more height out of this update?
- I need to follow up with my request for quote of an extension of ~10" height.
Mid afternoon I moved some equipment to QIL for making Q measurements of cantilevers, which Karthik is planning to do in the IR labs cryostat. See cryo elog for more information, photos attached to show location of equipment in QIL.
This morning I rolled the AG4395A from Adaptive Optics lab (labeled QIL, IP = 10.0.1.64) for use in Crackle. Will return after end using.
doesn't seem so, but they sell this one:
which has a USB interface and pretty good voltage noise spectrum
Is the reverse bias programmable? FEMTO has a bias trimmer on it. It's useful in the usual application, but for automation, the configuration of the input becomes cumbersome.
I was thinking about getting this new current pre-amp from NF:
It seems to have a good noise performance and has a built in low pass filter and also a remote interface.
The FEMTO seems less fancy, but their noise performance is actually 2-3x better.
FEMTO DLPCA200 low noise preamp (brand new)
Keithley Source Meter 2450 (brand new) => Returned 11/23/2020
were brought to the OMC lab for temporary use.
I have a python script for communication with the Met One 227a particle counter, but it appears like I am not receiving a response from the device. I swapped out a few serial-USB cables, but this did not fix the issue. I suspect I am making an unsound assumption about the communication protocol so I am going back to basics and reading more on RS232 and ASCII commands.
I moved the brand new TED200C on the workbench to Crackle for 2um ECDL (permanently)
The TED200C temp controller used in the 2um PD test setup will stay there (permanently)
Here's the python code I used to control this.
I incorrectly used the Move to Limit command ('1MV-3': axis 1, MoVe, negative direction, speed 3', where the speeds are given in the manual, see Section 4.7 in particular). Once this command is issued, the stage will keep moving until it receives the stop command. The JOG command would be more appropriate.
I confirmed a smooth change in the PD output as the beam translated across it.
I installed the Agilis mirror before the lens and cryo-chamber. Used the USB interface to align the beam onto the PD. So we can control the alignment remotely now (or once I’ve properly connected the USB cable instead of today’s janky test connection).
I installed the Agilis mirror before the lens and cryo-chamber. Used the USB interface to align the beam onto the PD. So we can control the alignment remotely now (or once I’ve properly connected the USB cable instead of today’s janky test connection).
An issue was raised with last calculation about the fact that our sensing of PDH signal isn't ideal and in the real world there is scattering, clipping extra adding excess noise in the PDH loop. This noise primarily comes by the intensity noise imparted on promptly reflected light from the cavity via various shaking optics etc on the table before it goes to the PDH reflection RF photodiode.
This noise's coupling to the PDH loop is identical to how shot noise of light couples into the PDH loop i.e.:
Looks like the temperature difference between the PD and the shield is relatively small. Even the transients when the heater is applied are order 5K.
This means that, for quick purposes, the shield RTD is a good proxy for the PD temperature.
The attached data is the difference between PD and shield RTD from circa 5th-6th February 2020.
Okay - all the steps in the procedure of eLOG 2476 have been verified as working - with the exception of the RTDs in the chamber.
With regards to taking dark noise spectra at different biases and temperatures, looks like Raymond took spectra with biases of [50, 100, 200, 400, 600, 1000]mV. If no objections, I’ll stick to that number of measurements.
I’m a bit pushed for time with other stuff. I wonder if the shield RTD is sufficient to run tests on the system? I’ll go back through the data and see how reproducible the relationship between shield temperature and PD temperature is. If it is reliable then in the interests of time, I’m going to forgo re-installing the extra RTDs in the chamber just now.
I have a preliminary calculation to post here. This does not include noise sources from cavity fluctuations and main frequency noise. But it gives some idea about shot noise and frequency noise of AUX laser conttribution to the noise in calibration.
Embellished Chris's PD MEDM screen a bit to illustrate controls in a diagram. The representation of the RELAY SWITCH between the Keithley and the SR560 is a bit off - I think the transimpedance amplifier is switched out as well.
Also - Keithley bright PD sweep output is attached.
Quick update, more detailed update to follow.
Still to do:
If we use ECDL for auxiliary frequency in 40m and hope to stabilize it up to 1 MHz with digital compensation of PZT, it is important to take into account any phase effect of the nearby FSR at 3.97 MHz. This should ideally be included in the Input Mode Cleaner loop considerations as well. These effects would be more prominent in longer cavities like aLIGO and LISA where FSR is very low and should we attempt to stabilize a laser lock beyond cavity's FSR.
I did a no assumptions calculation for getting a general transfer function fo PDH error signal in units of [W/Hz] assuming 1 W of incident power. This calculation would soon be uploaded here. I'll put down here primary results.
For incident field on a Fabry-Perot cavity (with fsr of ), reflected electric field transfer function (unitless) is given by:
Then, PDH error signal for a modulation frequency of at a modulation index of , in units of [W/Hz] (i.e. error signal power per Hz of error in laser frequency from cavity resonance) is given by:
after demodulation and low pass filtering. Note this transfer function is a complex quantity as it carries phase information of the transfer function too. The real signal is obtained by multiplying this signal at with and taking the real value of the product.
Having done this, we can see how in the real PDH error signal, there is a low pass at cavity pole, given by and a notch every fsr. The notch creates a zig-zag in the phase of the tranfer function and has a HWHM same as cavity pole. After this point, I just fitted a ZPK model to the transfer function to obtain a empirically derived model for PDH error signal transfer function. Apart from the cavity pole, this model needs to have resonance and antiresonance features present at each FSR with resonance having a linewidth of cavity pole while anti-resonance having a linewdth of . Here's how the ZPK model would look like:
I've attached my notebook where I did the fitting analysis and the overlap plot of real PDH error signal TF and the modelled approximation.
would be easier to achieve there with higher laser powers and higher cavity finesse.
But I haven't attempted that here as we do not know our NPRO PZT's resonance features yet.
I don't know why it would be easier to have higher finesse with longer arms. Something about beam size???
The NPRO PZT TF's are all in the 40m elog - there are many measurements of TF made over the past 10 years. Its like Raiders of the Lost Ark - you have to believe its there while searching.
Following up on the last post, here I presented a near back of the envelope calculation of how different choices of AUX cavity finesse and laser source for mariner would affect the prospects of calibration scheme.
As mentioned in the last elog post, here I considered using an NPRO seeded auxiliary laser source (converted to 1418nm by whatever method), ECDL based on ANU design with a modified PDH loop and same ECDl with a digital compensation of PZT resonances. I have taken the residual frequnecy noise of these lasers as the dominant noise source in the calibration scheme. Craig and Gautam in their proposal for SoCal wanted the AUX laser to be locked to the arm cavity in a PDH shot noise limited way. That would be necessary for 4km interferometers and would be easier to achieve there with higher laser powers and higher cavity finesse.
Here I considered three cases. First assumes about 3% transmittance of 1418nm in ITM and ETM HR coatings for mariner. This gives a finesse of about 100 and a cavity pole of 18.9 kHz. I believe this is the existing case at 40m. Next we consider transmittance of 0.5% and 0.05% (500 ppm) of 1418nm in ITM and ETM HR coatings for mairner. These cases give finesse of 625 and 6.28k respectively with cavity poles at 3 kHz and 299 Hz respectively.
Page 1: Consideres the case of finesse of 100. The green dashed line shows the amount of drive strength (in m) required at different frequencies if we use ECDL with PZT resonance compensation, to get an SNR of 1000 in 100s of integration time.
Page 2: Same as above but for Finesse of 625.
Page 3: Same as bove but for Finesse of 6280.
Page 4: Comparison of different finesse cases for the ECDL with PZT compensation option. Dashed curves represent requried drive strength (in m) for different cases.
Page 5: Same as above but for NPRO seeded auxiliary laser.
Note: For the NPRO seeded auxiliary laser, we have assumed that the noise of conversion to 1418 nm is similar to noise due to SHG process which is not dominant. There would be an effect of multiplying with a factor ranging form 1-1.5 due to frequency conversion but I have ignored it here for simiplicity. Also, NPRO case is limited in bandwidth due to PZT resonances. We might be able to get away with them using digital compensation like the case study for ECDL. But I haven't attempted that here as we do not know our NPRO PZT's resonance features yet.
We can use Thorlabs SAF1450S2 gain chip to generate 1418 nm light using an ECDL design similar to the one described in Kapasi et al. Optics Express Vol. 28, Issue 3, pp. 3280-3288 (2020) (ANU 2um ECDL design).
I have contacted Disha and Johannes to get the actual measured data for the PZT transfer function of this ECDL design. Fig.5b in their paper plots the transfer function of the PZT. Since, in ECDL PZT directly changes the cavity length, it has a more powerful actuation strength (2 orders of magnitude more) with actuation of 560 MHz.V upto 1 kHz. It however had a very low pole at 1 kHz and two mechanical resonance-antiresonance pairs near 1 kHz and 2 kHz. I modeled a transfer function by eye using Fig.5b of the paper. Page 1 in the attached pdf shows this modelled transfer function.
Next, we need to change the PDH loop for the auxiliary laser lock with the 40m cavity since the PZT has changed. I modelled one from scratch. This simple analog loop's performance is shown in orange in pages 2-5. This loop seemed stable from all the metrics I know, viz: phase margin of about 55 degrees (Page 2), no strong peak in close loop transfer function (page 3), and no remanant oscillations in time domain response (page 4).
I also modeled a similar loop but with digital compensation of the resonance-antiresonance features. This loop is plotted in green on pages 2-5. Both these loops have 300 kHz of bandwidth just by using PZT. I beleive this could be increased but I have not taken into account any saturation of PZT.
From Fig.4. of the paper gives a frequency noise estimate for free running ECDL. They mentioned that a roll-off below 10 Hz was due to their thermal feedback to remain in linear range of their frequency noise emasruement method. I modeled the noise of ECDL hence by
where the flicker noise contribution is similar to NPRO noise but ECDL has a white noise of 15 Hz/rtHz due to natural linewidth of spontaneous emission or Schawlow-Townes linewidth (with several broadening factors). I think this is an inherent limitation of ECDLs.
Page 5 shows both unsuppressed and suppressed frequency noise estimate for ECDL with the loops mentioned above and current values of NPRO noise are also plotted for comparison.
Came into lab today, with two main goals:
1) Bring Downs equipment for vibration measurements of cryo cooler during next operation
--> DONE, see photo
2) Assemble shields, with two layer scheme using electropolished and mill finish panels.
--> issue: the electropolished shield set was only partially shipped out in Ticket 15101 - unfortunately there was a misinterpretation by Logistics personnel regarding which parts to ship, due to the formatting of the ticket page. I will coordinate shipment of the balance of the panels to the vendor Able Electropolishing for completion of PO S479514